OpenStudy (anonymous):

HELLLLPPPPP!! A fair die has one face numbered 1,one face numbered 3,two faces numbered 5 and two faces numbered 6. find the probability of obtaining at least 7 odd numbers in 8 throws of the die.

OpenStudy (anonymous):

Pr( getting 7 odds in at least 8 throws) = 1 - Pr (getting 7odds in less than 8 throws = 1 - Pr (getting 7 odds in 7 throws) = 1 - (4/6)^7 = 1 - (2/3)^7 = 2059 / 2187

OpenStudy (anonymous):

And yes, that's a nasty looking fraction in the answer.. It's close to one as expected. Usually you're gonna have to throw it at least 8 times to get 7 odds unless you get really lucky and get 7 odds in a row to start.

OpenStudy (anonymous):

answer is .0195 so,above answer is wrong Ps:the question says 'AT LEAST' 7 odd numbers..it could be 8 odd numbers too

OpenStudy (anonymous):

The way I understand the question is that you must find the probability it will take at least 8 throws to get 7 odd. Under that interpretation, my answer is correct. I don't see any other way to interpret it. Maybe there is , but I don't see it.

OpenStudy (anonymous):

If 4 sides are odd and 2 even, it's usually going to take you at least 8 throws to get your 7 odds, so a small probability of .0195 makes no sense at all.

OpenStudy (anonymous):

No,,, you said 7 odds in at least 8 throws. did you mean at least 7 odds in 8 throws??

OpenStudy (anonymous):

that's a huuuuge difference!

OpenStudy (anonymous):

My bad. it's really 'at least 7 odds in 8 throws'.